tom is 10 years older than Kim. Next year, he will be twice as old as Kim. How old is Tom this year?
Answer by Maths68(1474)

Let
t = age of Tom
k=age of Kim
Tom is 10 years older than Kim means if we add 10 years to Kim age both will be of same age
t=k+10 …………..(1)
Next year means add 1 year to the age of kim and tom both
t+1=2(k+1)
t+1=2k+2
t=2k+2-1
t=2k+1……………(2)
Put the value of t from (2) to (1)
2K+1=k+10
2k=k+10-1
2k=k+9
2k-k=9
k=9
Put the value of k in (1)
t=k+10
t=9+10
t=19
Tom is 19 years old and Kim is 9 years old.